The net income of Novis Corporation is $95,000. The company has 20,000 outstanding shares and a 100 percent payout policy. The expected value of the firm one year from now is $1,805,000. The appropriate discount rate for the company is 12 percent, and the dividend tax rate is zero. a. What is the current value of the firm assuming the current dividend has not yet been paid? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Current value of the firm $ b. What is the ex-dividend price of the company’s stock if the board follows its current policy? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Ex-dividend price $ c. At the dividend declaration meeting, several board members claimed that the dividend is too meager and is probably depressing the company’s price. They proposed that the company sell enough new shares to finance a $5.95 dividend. c-1. Calculate the current value of the firm. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Current value of the firm $ c-2. If the proposal is adopted, at what price will the new shares sell? How many will be sold? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.) Current share price $ Shares sold
Answer #1

Given get income = $95000 no ef shares 20000 Payent Policy & 100% other one year value of fish $1808ooa) Disco Pak = 12% (ke)

ff Partb cassent value of fish (Po) dividend not get paid dos P4 & It ke Po = 1805060 - = $1611607.14 17.72 per Shave a 80.86

fast D grow value of the finu Cas for mm Apprally npo - holt (htm) Pg - mei itke no no. of shore at begining of yeary me no.

= 2000045.95+ (200004 265.93) 90.85 (285-93 r qu257 I foye = 119000 t 1829000 - 24000 112 1924ooo هر | Inpo = 17 17857.ly 1 C